# A Survey on Data Selection for Language Models

### v2

@@ 9 Related Topics
There are a number of data-focused topics in machine learning that are related to data selection, but have separate aims and use cases.
@@
Dataset distillation (also known as dataset condensation) and coreset selection (also known as data pruning) have been extensively studied for reducing the size of training data. Dataset distillation aims to synthesize a small dataset such that the model trained on that synthetic dataset achieve performance akin to one trained on the full dataset. Dataset distillation is first introduced by Wang et al. (2018b), where the model weights are treated as a function of the distilled dataset (outer optimization) and the distilled dataset is optimized (inner optimization) so that randomly initialized models trained on it with one-step gradient descent could achieve good performance. Such a bi-level meta-learning framework often incurs substantial computational costs. To improve the scalability and effectiveness of dataset distillation, subsequent research (Zhao et al., 2020; Lee et al., 2022b; Zhao & Bilen, 2021) has employed gradient matching as the supervision signal, aligning the gradients of model parameters trained on the original dataset and the distilled dataset. Moreover, researchers have proposed techniques such as feature distribution matching (Wang et al., 2022a; Zhao & Bilen, 2023; Zhao et al., 2023), training trajectory matching (Cazenavette et al., 2022; Cui et al., 2023b; Guo et al., 2024), and kernel methods  (Nguyen et al., 2020; 2021; Zhou et al., 2022).
@@
Data valuation and attribution focus on understanding how each training data point affects a models predictions. Methods for valuation and attribution can help to identify potentially mislabeled data, identify brittle predictions, quantify the affect of a training data point on an evaluation data point, and even quantify train-test contamination. These methods tend to be computationally expensive, although recent progress has improved their efficiency.
In recent work, Ghorbani & Zou (2019) propose Data Shapley, a principled framework for data valuation, where the goal is to quantify the value of each data point to a predictors performance. The Data Shapley value for each data point can then be used as the utility function in a data selection method for either distribution matching or diversification purposes. Ilyas et al. (2022) present datamodels, a framework used to predict a models performance on a data point $x^{(i)}$, given that the model is trained on a known subset of training data. They demonstrate that even using a linear function as the predictor, their datamodelling framework is capable of accurately predicting model outputs and counterfactuals, as well as encoding a metric for data point similarity. Similarly, Park et al. (2023) propose TRAK, a data attribution method that can make accurate counterfactual predictions, trying to answer the question “why did my model make this prediction”. They show that TRAK is a plausible method for determining which training data was most impactful on a models prediction.
@@
Data augmentation methods are used to generate new data by modifying the existing data points in a dataset. Similar to data selection, data augmentation also aims to shift the training distribution. Data selection is often used to improve the data coverage in domains where there is limited data (Li et al., 2022b; Chen et al., 2023a) and can be used in conjunction with existing data selection methods (e.g. domain-specific selection in §3.4 or “selecting from auxiliary data to improve generalization” in §7). For example, Albalak et al. (2022a) use a policy gradient along with information from an auxiliary task to improve their main model. However, the quality of the coverage produced from augmentation may be limited, as the semantics of augmented data can diverge from the underlying distribution of language, leading to an undesirable distribution shift and possibly a decrease in performance (Li et al., 2022b; Chen et al., 2023a). Thus, care should be taken when performing data augmentation to ensure that the augmentations do not stray too far from the desired distribution. Nonetheless, data augmentation can still be a useful tool, and we refer readers to existing surveys on the many methods for augmentation in NLP (Chen et al., 2023a; Feng et al., 2021), code (Zhuo et al., 2023), and computer vision (Yang et al., 2022).
@@
Data curation is a set of processes surrounding the discovery, organization, integration, annotation, cleaning, storage and maintenance of data (McLure et al., 2014; Freitas & Curry, 2016; Thirumuruganathan et al., 2020). Some of the methods that are used for curation overlap with those from data selection (e.g. cleaning and filtering), but there are also methods from data selection which are not included in data curation (e.g. heuristic filters, data mixing, and filtering toxic and explicit content). Data selection is, generally speaking, a process that occurs after data curation, once a practitioner is prepared to train a model. After the data has been curated, then data selection methods can be used to create a dataset for the desired purpose.
@@
While data selection aims to determine which data points a model should train or evaluate on, and how many times they may be repeated for training, curriculum learning aims to improve a model through an orthogonal goal. Curriculum learning is often motivated by the idea that the training data should be increasing in complexity according to the models current capability, and aims to determine when a data point should be trained on.
Curriculum learning has found limited success in the natural language domain. For example, Graves et al. (2017) use multi-armed bandits to determine an appropriate curriculum for language modeling with LSTMs. Xu et al. (2020) demonstrate that curricula can be used to improve learning in natural language understanding tasks for a BERT model. More recently, Fan & Jaggi (2023) propose irreducible curriculum for language model pretraining, which prioritizes training on samples with higher learnability (motivated by RHO-Loss (Mindermann et al., 2022)). Curriculum learning is one aspect of data-centric AI that has not yet seen a large adoption in large language models, in part due to its limited success.
@@ 10.1 Test set decontamination
Understanding a model’s capabilities requires evaluation on a fair benchmark, but the training of language models on web-scale data has raised concerns and speculation about the possible inclusion of evaluation benchmarks in training data. Given the scale of data it is not feasible to manually check, so automated methods have been designed to detect the existence of evaluation data in the training set (Rae et al., 2021; Li et al., 2022a; 2023b; Magnusson et al., 2023; Elazar et al., 2023), as well as detecting models that have already been trained on evaluation data (Carlini et al., 2018; Jagielski, 2023; Marone & Van Durme, 2024). Jacovi et al. (2023) on the other hand, suggest three strategies for reducing the risks of contamination. Data decontamination is an important step in ensuring the integrity of model evaluation, and is included in the discussion section because this critical step is relevant, and should be required, for all stages of model training.
@@
Test set contamination methods often follow a similar form to deduplication methods. When decontaminating MassiveTextRae et al. (2021) compute the 13-gram Jaccard similarity between train and test documents (as originally proposed by Lee et al. (2022a)) and remove train documents that have a Jaccard similarity over 0.8, the same exact method that they use to detect duplicated documents. Similarly, Brown et al. (2020) and Gao et al. (2020) use 13-gram overlap filtering to detect contamination in their respective training sets. To make decontamination easier, some tools have been developed. For example, CarperAI create a GitHub repository with code for decontaminating evaluation data from a training dataset using MinHashLSH. [^20] CarperAI decontamination GitHub repository Additionally, Marone & Van Durme (2024) propose Data Portraits, a method of investigating the data that a model was trained on. Data portraits are an artifact that records training data, allows downstream inspection, and enables answering questions about test set leakage and model plagiarism through a novel membership testing tool, a strided bloom filter. In addition to the new method, they release a demo of their tool [^21] https://dataportraits.org/ which uses 50-character grams to detect overlap between training data and an input sequence.
@@ 10.2 Tradeoffs between memorization and generalization
While most methods of data selection discussed here, data deduplication in particular, suggest that memorization is a negative model property, this is not always the case. There are truths about the world which always hold and memorization is indeed a desirable property in those cases. For example, when answering the question “what is the formula for Pythagoras’ theorem?” we expect a model’s answer to deviate very minimally from “$a^{2}+b^{2}=c^{2}$” and a serious deviation can have real world consequences if such a model was used in an education setting. However, generalization is still desirable in this situation as changing the variable names should still be recognized as correct (e.g.$s_{1}^{2}+s_{2}^{2}=s_{3}^{2}$”). Furthermore, memorization is also desirable for code data, such as specific API calls in tool usage for LLMs (Schick et al., 2023; Pan et al., 2023), and syntax for coding languages (Chen et al., 2021; Li et al., 2022a; Allal et al., 2023). Memorization is desirable in many other contexts as well, and Brown et al. (2021) have theoretically proven that in many cases memorization is a requirement of any reasonably accurate training algorithm.
On the other hand, it has been demonstrated that not only do models memorize the desirable information, but also the undesirable information such as personally identifiable information, phone numbers, credit card information and much more. Carlini et al. (2020) demonstrate this with a simple attack for extracting verbatim sequences of an LLMs training data, finding that although data points from the training set do not have noticeably lower loss than test data points on average, there are some individual examples that are indeed memorized.  Biderman et al. (2023a) study whether it is possible to predict which sequences from the training data will be memorized by using smaller models, or partially trained models.  Carlini et al. (2023) demonstrate that memorization grows significantly with three factors: the capacity of a model, the number of times a sample is repeated in the training data, and the number of tokens used to prompt the model to emit the memorized data. While the third factor, the number of tokens in a prompt, is not in the control of model developers, the first two are. Model developers can choose to use smaller models to satisfy the needs for a given use case, rather than relying on massive models which generalize across many use cases, but tend to memorize data. Furthermore, deduplication has been demonstrated as a strong method for minimizing memorization. However, as previously discussed, deduplication needs to be intelligently applied so that “good” memorization can occur (e.g. facts), while reducing “bad” memorization (e.g. PII).
@@ 10.3 There’s no free lunch for filtering
Overly filtering the data can lead to undesirable biases in the model. In particular, Dodge et al. (2021) find that blocklist filtering used to create C4 (Raffel et al., 2020) disproportionately removes text written by, and about, minority individuals. Additionally, Welbl et al. (2021) find that removing toxic content (as determined by the Jigsaw Perspective API [^23] https://perspectiveapi.com/ can lead to worse evaluation loss and disproportionately affects texts about and by marginalized groups. This finding was also replicated by Longpre et al. (2023c). Additionally, in spite of the filtering performed in C4Luccioni & Viviano (2021) still find significant quantities of sexually explicit content and hate speech. These works demonstrate the need for filtering strategies with improved precision (remove less text that is desirable) and recall (remove more text that is undesirable).
@@ 10.4 Tools for data selection
When performing data selection, a first step should generally be to understand the raw dataset, and a number of tools have been developed to assist in data exploration and auditing. For example, researchers at AI2 developed a search engine over the C4 dataset, [^24] https://c4-search.apps.allenai.org/ and Piktus et al. (2023) create a search engine over C4, the Pile, ROOTS, and text captions from LAION.252525code at https://github.com/huggingface/gaia, demonstration at https://huggingface.co/spaces/spacerini/gaia is not functioning as of 02/07/2024  Elazar et al. (2023) describe their What’s in my big data tool, which can search for and count n-grams in a number of popular pretraining corpora. Additionally, their online demo contains tools for exploring the internet domains included in datasets, a visualization of the amount of text overlap between corpora, and a visualization to understand the length of documents in each corpus. [^26] https://wimbd.apps.allenai.org/ Finally, Liu et al. (2024a) developed the Infini-gram which uses suffix arrays as a method for language modeling, and provide a demonstration of the Infini-gram trained on the RedPajama, the Pile, and Dolma corpora and is capable of very quickly searching for n-grams within any of those corpora.  [^27] https://huggingface.co/spaces/liujch1998/infini-gram
@@
Once the data has been explored and audited, the next step of data selection is to devise the appropriate selection mechanisms and apply them over the data. Some of the methods discussed in this work have been implemented as open-sourced tools, allowing anyone to easily make use of them. The CCNet pipeline (Wenzek et al., 2020)  [^28] https://github.com/facebookresearch/cc_net is a commonly used tool for downloading and cleaning Common Crawl data. The CCNet pipeline performs deduplication at the line level, performs language identification with a fastText linear classifier, and can train an n-gram language model to calculate perplexity per document. Soldaini et al. (2024) open-source tools for identifying non-language text, language identification, personally identifiable information, toxic text, text quality, and duplicated data. [^29] https://github.com/allenai/dolma Xie et al. (2023b) provide a pip-installable tool for selecting data with a distribution similar to a target dataset, that can be used either for domain-specific selection, or for quality filtering. [^30] https://github.com/p-lambda/dsir Lee et al. (2022a) provide the code to run their ExactSubstr deduplication, [^31] https://github.com/google-research/deduplicate-text-datasets and Computer (2023) provide tools for computing exact duplicates with a bloom filter, and fuzzy deduplication with locality sensitive hashing. [^32] https://github.com/togethercomputer/RedPajama-Data Additionally, the tools provided by Computer (2023) contain code for calculating a wide variety of data signals including heuristics, quality, duplication, and toxicity. Finally, CarperAI provide a pip-installable package for decontaminating datasets from a number of openly available benchmarks. [^33] https://github.com/CarperAI/decontamination/tree/main

### v3

@@ 9 Related Topics
Dataset distillation (also known as dataset condensation) and coreset selection (also known as data pruning) have been extensively studied for reducing the size of training data. Dataset distillation aims to synthesize a small dataset such that the model trained on that synthetic dataset achieve performance akin to one trained on the full dataset. Dataset distillation is first introduced by Wang et al. (2018b), where the model weights are treated as a function of the distilled dataset (outer optimization) and the distilled dataset is optimized (inner optimization) so that randomly initialized models trained on it with one-step gradient descent could achieve good performance. Such a bi-level meta-learning framework often incurs substantial computational costs. To improve the scalability and effectiveness of dataset distillation, subsequent research (Zhao et al., 2020; Lee et al., 2022b; Zhao & Bilen, 2021) has employed gradient matching as the supervision signal, aligning the gradients of model parameters trained on the original dataset and the distilled dataset. Moreover, researchers have proposed techniques such as feature distribution matching (Wang et al., 2022a; Zhao & Bilen, 2023; Zhao et al., 2023a), training trajectory matching (Cazenavette et al., 2022; Cui et al., 2023b; Guo et al., 2024), and kernel methods  (Nguyen et al., 2020; 2021; Zhou et al., 2022).
@@
Data valuation and attribution focus on understanding how each training data point affects a model’s predictions. Methods for valuation and attribution can help to identify potentially mislabeled data, identify brittle predictions, quantify the affect of a training data point on an evaluation data point, and even quantify train-test contamination. These methods tend to be computationally expensive, although recent progress has improved their efficiency.
In recent work, Ghorbani & Zou (2019) propose Data Shapley, a principled framework for data valuation, where the goal is to quantify the value of each data point to a predictors performance. The Data Shapley value for each data point can then be used as the utility function in a data selection method for either distribution matching or diversification purposes. Ilyas et al. (2022) present datamodels, a framework used to predict a models performance on a data point $x^{(i)}$, given that the model is trained on a known subset of training data. They demonstrate that even using a linear function as the predictor, their datamodelling framework is capable of accurately predicting model outputs and counterfactuals, as well as encoding a metric for data point similarity. Similarly, Park et al. (2023) propose TRAK, a data attribution method that can make accurate counterfactual predictions, trying to answer the question “why did my model make this prediction”. They show that TRAK is a plausible method for determining which training data was most impactful on a model’s prediction.
@@
Data augmentation methods are used to generate new data by modifying the existing data points in a dataset. Similar to data selection, data augmentation also aims to shift the training distribution. Data selection is often used to improve the data coverage in domains where there is limited data (Li et al., 2022b; Chen et al., 2023a) and can be used in conjunction with existing data selection methods (e.g., domain-specific selection in §3.4 or “selecting from auxiliary data to improve generalization” in §7). For example, Albalak et al. (2022a) use a policy gradient along with information from an auxiliary task to improve their main model. However, the quality of the coverage produced from augmentation may be limited, as the semantics of augmented data can diverge from the underlying distribution of language, leading to an undesirable distribution shift and possibly a decrease in performance (Li et al., 2022b; Chen et al., 2023a). Thus, care should be taken when performing data augmentation to ensure that the augmentations do not stray too far from the desired distribution. Nonetheless, data augmentation can still be a useful tool, and we refer readers to existing surveys on the many methods for augmentation in NLP (Chen et al., 2023a; Feng et al., 2021), code (Zhuo et al., 2023), and computer vision (Yang et al., 2022).
@@
Data curation is a set of processes surrounding the discovery, organization, integration, annotation, cleaning, storage and maintenance of data (McLure et al., 2014; Freitas & Curry, 2016; Thirumuruganathan et al., 2020). Some of the methods that are used for curation overlap with those from data selection (e.g., cleaning and filtering), but there are also methods from data selection which are not included in data curation (e.g., heuristic filters, data mixing, and filtering toxic and explicit content). Data selection is, generally speaking, a process that occurs after data curation, once a practitioner is prepared to train a model. After the data has been curated, then data selection methods can be used to create a dataset for the desired purpose.
@@
While data selection aims to determine which data points a model should train or evaluate on, and how many times they may be repeated for training, curriculum learning aims to improve a model through an orthogonal goal. Curriculum learning is often motivated by the idea that the training data should be increasing in complexity according to the models current capability, and aims to determine when a data point should be trained on. Curriculum learning has found limited adoption in the natural language domain. For example, Graves et al. (2017) use multi-armed bandits to determine an appropriate curriculum for language modeling with LSTMs. Xu et al. (2020) demonstrate that curricula can be used to improve learning in natural language understanding tasks for a BERT model. More recently, Fan & Jaggi (2023) propose irreducible curriculum for language model pretraining, which prioritizes training on samples with higher learnability (motivated by RHO-Loss (Mindermann et al., 2022)). Curriculum learning is one aspect of data-centric AI that has not yet seen a large adoption in large language models, in part due to its limited success.
Active learning.
Active learning is a long-studied problem that aims to answer the question: given unlabeled data points, how should a fixed annotation budget be used to optimally train a model for supervised tasks, and has been applied to a wide variety of settings ranging from text classification (Lewis & Gale, 1994) to in-context learning (Wu et al., 2023). Active learning is closely related to data selection as both fields consider the question of how to find the optimal data to train a model, but focus on different settings. Unlike data selection, active learning explicitly focuses on supervised training. Please see the available surveys on active learning in NLP (Zhang et al., 2022b), and general deep learning (Ren et al., 2021; Zhan et al., 2022; Tharwat & Schenck, 2023) for a deeper discussion on the topic.
@@ 10.1 Test set decontamination
Understanding a model’s capabilities requires evaluation on a fair benchmark, but the training of language models on web-scale data has raised concerns and speculation about the possible inclusion of evaluation benchmarks in training data. Given the scale of data it is not feasible to manually check, so automated methods have been designed to detect the existence of evaluation data in the training set (Rae et al., 2021; Li et al., 2022a; 2023b; Magnusson et al., 2023; Elazar et al., 2024), as well as detecting models that have already been trained on evaluation data (Carlini et al., 2018; Jagielski, 2023; Marone & Van Durme, 2024). Jacovi et al. (2023) on the other hand, suggest three strategies for reducing the risks of contamination. Data decontamination is an important step in ensuring the integrity of model evaluation, and is included in the discussion section because this critical step is relevant, and should be required, for all stages of model training.
@@
Test set contamination methods often follow a similar form to deduplication methods. When decontaminating MassiveTextRae et al. (2021) compute the 13-gram Jaccard similarity between train and test documents (as originally proposed by Lee et al. (2022a)) and remove train documents that have a Jaccard similarity over 0.8, the same exact method that they use to detect duplicated documents. Similarly, Brown et al. (2020a) and Gao et al. (2020) use 13-gram overlap filtering to detect contamination in their respective training sets. To make decontamination easier, some tools have been developed. For example, CarperAI released software for decontaminating evaluation data from a training dataset using MinHashLSH. [^21] CarperAI decontamination GitHub repository Additionally, Marone & Van Durme (2024) propose Data Portraits, a method of investigating the data that a model was trained on. Data portraits are an artifact that records training data, allows downstream inspection, and enables answering questions about test set leakage and model plagiarism through a novel membership testing tool, a strided bloom filter. In addition to the new method, they release a demo of their tool [^22] https://dataportraits.org/ which uses 50-character grams to detect overlap between training data and an input sequence. Elazar et al. (2024) indexed several corpora using elastic-search, a reverse index, and used it to detect exact matches of multiple test-sets used for evaluation. Finally, Merrill et al. (2024) built an efficient automtaton named CDAWG (Inenaga et al., 2005), that allows a flexible n-gram match to the longest suffix that matches an input string in the (indexed) training data.
@@ 10.2 Tradeoffs between memorization and generalization
While most methods of data selection discussed here, data deduplication in particular, suggest that memorization is a negative model property, this is not always the case. There are truths about the world which always hold and memorization is indeed a desirable property in those cases. For example, when answering the question “what is the formula for Pythagoras’ theorem?” we expect a model’s answer to deviate very minimally from “$a^{2}+b^{2}=c^{2}$” and a serious deviation can have real world consequences if such a model was used in an education setting. However, generalization is still desirable in this situation as changing the variable names should still be recognized as correct (e.g., “$s_{1}^{2}+s_{2}^{2}=s_{3}^{2}$”). Furthermore, memorization is also desirable for code data, such as specific API calls in tool usage for LLMs (Schick et al., 2023; Pan et al., 2023), and syntax for coding languages (Chen et al., 2021; Li et al., 2022a; Allal et al., 2023). Memorization is desirable in many other contexts as well, and Brown et al. (2021) have theoretically proven that in many cases memorization is a requirement of any reasonably accurate training algorithm.
On the other hand, it has been demonstrated that not only do models memorize the desirable information, but also the undesirable information such as personally identifiable information, phone numbers, credit card information and much more. Carlini et al. (2020) demonstrate this with a simple attack for extracting verbatim sequences of an LLMs training data, finding that although data points from the training set do not have noticeably lower loss than test data points on average, there are some individual examples that are indeed memorized.  Biderman et al. (2023a) study whether it is possible to predict which sequences from the training data will be memorized by using smaller models, or partially trained models.  Carlini et al. (2023) demonstrate that memorization grows significantly with three factors: the capacity of a model, the number of times a sample is repeated in the training data, and the number of tokens used to prompt the model to emit the memorized data. While the third factor, the number of tokens in a prompt, is not in the control of model developers, the first two are. Model developers can choose to use smaller models to satisfy the needs for a given use case, rather than relying on massive models which generalize across many use cases, but tend to memorize data. Furthermore, deduplication has been demonstrated as a strong method for minimizing memorization. However, as previously discussed, deduplication needs to be intelligently applied so that “good” memorization can occur (e.g., facts), while reducing “bad” memorization (e.g., PII).
@@ 10.3 There’s no free lunch for filtering
Overly filtering the data can lead to undesirable biases in the model. In particular, Dodge et al. (2021) find that blocklist filtering used to create C4 (Raffel et al., 2020) disproportionately removes text written by, and about, minority individuals. Additionally, Welbl et al. (2021) find that removing toxic content (as determined by the Jigsaw Perspective API [^24] https://perspectiveapi.com/) can lead to worse evaluation loss and disproportionately affects texts about and by marginalized groups. This finding was also replicated by Longpre et al. (2023c). These works demonstrate the need for filtering strategies with improved precision (remove less text that is desirable) and recall (remove more text that is undesirable).
@@ 10.4 Tools for data selection
When performing data selection, a first step should generally be to understand the raw dataset, and a number of tools have been developed to assist in data exploration and auditing. For example, researchers at AI2 developed a search engine over the C4 dataset, [^25] https://c4-search.apps.allenai.org/ and Piktus et al. (2023) create a search engine over C4, the Pile, ROOTS, and text captions from LAION.262626code at https://github.com/huggingface/gaia, demonstration at https://huggingface.co/spaces/spacerini/gaia is not functioning as of 02/07/2024  Elazar et al. (2024) describe their What’s in my big data tool, which can search for and count n-grams in a number of popular pretraining corpora. Additionally, their online demo contains tools for exploring the internet domains included in datasets, a visualization of the amount of text overlap between corpora, and a visualization to understand the length of documents in each corpus. [^27] https://wimbd.apps.allenai.org/ Finally, Liu et al. (2024a) developed the Infini-gram which uses suffix arrays as a method for language modeling, ultimately demonstrating competitive performance on the RedPajama, the Pile, and Dolma corpora in addition to being able to quickly search for n-grams within any of those corpora.  [^28] https://huggingface.co/spaces/liujch1998/infini-gram
@@
Once the data has been explored and audited, the next step of data selection is to devise the appropriate selection mechanisms and apply them over the data. Some of the methods discussed in this work have been implemented as open-sourced tools, allowing anyone to easily make use of them. The CCNet pipeline (Wenzek et al., 2020)  [^29] https://github.com/facebookresearch/cc_net is a commonly used tool for downloading and cleaning Common Crawl data. The CCNet pipeline performs deduplication at the line level, performs language identification with a fastText linear classifier, and can train an n-gram language model to calculate perplexity per document. Soldaini et al. (2024) open-source tools for identifying non-language text, language identification, personally identifiable information, toxic text, text quality, and duplicated data. [^30] https://github.com/allenai/dolma Xie et al. (2023b) provide a pip-installable tool for selecting data with a distribution similar to a target dataset, that can be used either for domain-specific selection, or for quality filtering. [^31] https://github.com/p-lambda/dsir Lee et al. (2022a) provide the code to run their ExactSubstr deduplication, [^32] https://github.com/google-research/deduplicate-text-datasets and Computer (2023) provide tools for computing exact duplicates with a bloom filter, and fuzzy deduplication with locality sensitive hashing. [^33] https://github.com/togethercomputer/RedPajama-Data Additionally, the tools provided by Computer (2023) contain code for calculating a wide variety of data signals including heuristics, quality, duplication, and toxicity. Additionally, CarperAI provide a pip-installable package for decontaminating datasets from a number of openly available benchmarks. [^34] https://github.com/CarperAI/decontamination/tree/main Finally, Penedo et al. (2024) provide a platform-agnostic python library that can filter and deduplicate data efficiently.